Aug 7, 2012 Readout

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • ESNET (Number is 8542553) and EVO session (GlueX Calorimetry meeting room)
  • Phone connection only upon request.
    • +1-866-740-1260 : US and Canada
    • +1-303-248-0285 : International
    • then use participant code: 3421244# (the # is needed when using the phone)
    • OR from type access code 3421244 into "join a meeting" without the # (you need java plugin)

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6337 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Items for followup from previous meeting

Presentations from recent meetings

Tentative Agenda

  1. Announcements
  2. Review of Action Items
  3. Update from Athens: monitoring
    • Initial version of support
  4. Update from USM
    1. SiPM Update (Javiera)
    2. Operational Voltage (Orlando)
    3. Light guides (Iñaki)
    4. Impacts of pixelation and cross-talk on signal linearity (Sergey)
  5. Status of testing at JLab
    1. Hall-D ELOG.
    2. Update on the analysis of the production units (Ben St. Laurent)
    3. Optimal bias voltage (Yi Qiang)
  6. Beam Tests, Cosmics, Simulations
    1. BCAL Beam Tests 2012
    2. BCAL hdgeant (David L.)
    3. BCAL Cosmics at Regina
    4. Simulations
  7. Other


Attending: Yi (JLab), Will, Sergey, Javiera, Orlando, Inaki, Hayk (USM), Andrei, Zisis (Regina)

  1. Announcements: None.
  2. Review of Action Items: None.
  3. Update from Athens: No report.
  4. Update from USM
    1. SiPM Update (Javiera): Work is progressing well. Next shipment of 600 units is in transit (Brazil) and should arrive in the USA on Thursday.
    2. Operational Voltage (Orlando): Correct equation 3 on p5, according to the discussion with Elton and Yi from two weeks ago. See plots starting on p34. Repeated measurements 70 times (dark rate as well). Correction to amplitude for cross talk was discussed.
    3. Light guides (Iñaki): Summary of calibration/Quality Control procedure on the machining was presented, in advance of the production process. The calibration of the machine should finish this week. In dimensional control the JLab gauge can be used only after polishing, as the machined light guides are now made 150 um oversize. Two sets (80 units) should be ready by the end of next week, and 680 units in a month or so. Question by Will: are these delays putting us on the critical path to cause delay in the project?
    4. Impacts of pixelation and cross-talk on signal linearity (Sergey): Sergey presented his simulation model on the effect of cross talk to Vop. He discussed the possible effect on energy resolution and compensation/correction ideas. Calibration of the SiPMs using hadrons (with help from CDC) could be used to check these issues.
  5. Status of testing at JLab: No report
  6. Beam Tests, Cosmics, Simulations
    1. BCAL Cosmics at Regina: Cosmics runs continue at 0.7 V overbias.
    2. Simulations: Sampling Fraction work continues. The extraction of the SF for each readout cell is stable now, and low energy effects are being examined using both GEANT3 and FLUKA.
  7. Other: nothing